Coal gasification is a process that converts
coal into methane gas.
2 C(s) + 2 H2O(ℓ) −→ CH4(g) + CO2(g)
If this reaction has a percent yield of 94.6%,
how much methane can be obtained from
5000 g of carbon?
1
Expert's answer
2017-03-31T13:08:07-0400
m teor (CH 4 ) = n(CH 4 )*M(CH 4 ) = n(C)/2*M(CH 4 ) = m(C)/M(C)/2*M(CH 4 ) = 5000/12/2*16 = 3333(g)
m(CH 4 ) = η*m teor (CH 4 ) = 0.946*3333 = 3153 (g)
Answer: 3153g methane can be obtained from 5000 g of carbon